Lead has a specific heat of 0.031 cal/gram °C. How many calories does it take to raise the temperature of 200 g of lead from 50°C to 80°C?
1
Expert's answer
2012-04-10T07:18:34-0400
The change of temperature is

& Delta;T = 80 - 50 = 30°C.

So it takes

& Delta;Q = 0.031cal/(gram °C) * 200g * 30°C = 186 cal

to raise the temperature of 200 g of lead from 50°C to 80°C.